Jo is 2 years older than Harry.
Kate is twice as old as Jo.
write an expression for each person's age using n
what is kate's age?

Answer by Alan3354(69443)

Jo is 2 years older than Harry.
Kate is twice as old as Jo.
write an expression for each person's age using n
---------------
n = Harry's age
Jo's age = n+2
Kate's = 2(n+2) = 2n+4
=========================
or,
n = Kate's age
Jo's age = n/2
Harry's age = (n/2) - 2 = (n-4)/2
======================
You can have Jo's age = n, too